# Diffusion approximation of the stationary distribution of a two-level single server queue

### v4

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study the limit of the stationary distribution of this two-level queue under a certain scaling in heavy traffic, which corresponds to a diffusion approximation. It is shown that this limiting distribution is truncated exponential below the threshold level and exponential above it under suitably chosen system parameters. This result is proved using the so called BAR approach studied in [2, 3, 10, 11]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distributions.
@@ 1 Introduction
This 2-level $GI/G/1$ queue is motivated by an energy saving problem to control the arrival process and service rate. Here, the system is an internet server, and its energy consumption depends on service rate. Then, the problem is to find an optimal level to minimize the operating energy of the server and customer’s waiting cost.
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. It may be considered to approximate using a quasi-birth-and-death process, and it is not hard to numerically compute its performance. However, it is hard to derive it in an analytically tractable form except for special cases such as all the distributions to be exponential. Because of this, a simple approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution. This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
@@
As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue (e.g., see Kingman [7]). Furthermore, Miyazawa [11] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room. Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see Harrison [5]). Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it. This conjecture is easily verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional. Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [2, 3, 10, 11]. This verification, Theorem 3.1, requires a certain limitation on system parameters or the exponentially distributed workload brought by a customer. The proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
@@
This paper is made up by five sections. We first introduce a general framework for a Markov process describing the 2-level $GI/G/1$ queue in Section 2, then provide notations and assumptions for a sequence of the 2-level $GI/G/1$ queues in heavy traffic in Section 3.1. A main result, Theorem 3.1, and the exponentially distributed case are presented in Section 3.2. In Section 4, we prove Theorem 3.1. Finally, in Section 5, we intuitively discuss two topics. In Section 5.1, we consider a process-limit for the 2-level $GI/G/1$ queue in heavy traffic. In Section 5.2, we intuitively verify the conjecture in Remark 3.1.
@@ 3.1 Assumptions
For the $r$-th model, let $\ell^{(r)}_{0}$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$. We refer to this number as the queue length. Let $t^{(r)}_{e,n}$ be the $n$-th arrival time of a customer, and let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customer. We assume that $T_{e}^{(r)}(n)\equiv t^{(r)}_{e,n}-t^{(r)}_{e,n-1}$ is independent of $\{t^{(r)}_{e,n^{\prime}};0\leq n^{\prime}\leq n-1\}$ and everything else for all $n\geq 1$, where $t^{(r)}_{e,0}=0$. Thus, the inter-arrival times are independent, but they have may two different distributions depending on the queue length. Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are random variables which may have different distributions. As for service, let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customers. We assume that $\{T^{(r)}_{s}(n);n\geq 1\}$ is a sequence of $i.i.d.$ random variable. We denote a random variable subject to the common distribution of $T^{(r)}_{s}(n)$ by $T^{(r)}_{s}$. Let $c^{(r)}_{1}$ and $c^{(r)}_{2}$ be the service speeds when the queue length is not greater than $\ell^{(r)}_{0}$ and otherwise, respectively. Namely, as long as $R^{(r)}_{s}(t)$ is differentiable at time $t$,
@@
and denote the counting processes of the arrivals and service completions by $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$, respectively. Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov. Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with respect to the time-shift operator $\Psi$. Let $b_{1}\in\mathbb{R}$ and $b_{2},\ell_{0}>0$ be constants, and assume the following assumptions.
@@
    $c^{(r)}_{i}\mu^{(r)}-\lambda^{(r)}_{i}=r\mu^{(r)}b_{i}+o(r)$ as $r\downarrow 0$, and $b_{2}>0$.
@@
    $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$, and $\ell^{(r)}_{0}\geq 2$.
@@
Note that (3.a) implies that
@@
Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$. Obviously, $\mathbb{E}[N_{e}(t)]\leq\mathbb{E}[N_{e,0}(t)]$ under the stationary framework. Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N_{e}(1)]\leq\mathbb{E}[N_{e,0}(1)]=1/\mathbb{E}[T_{e,1}(n)\wedge T_{e,2}(n)]<\infty$. Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)]<\infty$. ∎
@@ 3.2 Main results
We are now ready to present a main result, which will be proved in Section 4. Recall that $L^{(r)}$ is a random variable subject to the stationary distribution $\nu^{(r)}$ of the $r$-th 2-level $GI/G/1$ queue. We decompose $\nu^{(r)}$ into marginal measures $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$ as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
@@
$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}>\ell^{(r)}_{0}),$
@@
For each $r\in(0,1]$, define $\widetilde{\nu}_{i|cd}$ as the probability measures on ${\mathcal{B}}(\mathbb{R}_{+})$ which have the density functions $h_{i|cd}$ for $i=1,2$ which are
@@
Conversely, if (3.11) holds, then we have the weak convergence (3.10) from (i) because $\widetilde{A}_{i}>0$. Thus, (3.11) is equivalent to (3.10). We conjecture to have (3.10) without any condition in (ii). There are two supports for this conjecture. One is from a process limit for the 2-level $GI/G/1$ queue in heavy traffic. Another is from the intuitive verification of (3.11). They are detailed in Sections 5.1 and 5.2, repectively.
@@
Let consider Theorem 3.1 when the inter-arrival times and service workload are all exponentially distributed. We call this case exponential.
@@
Under the settings of Theorem 3.1, if $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed, then, for any $\lambda_{1}$ and $\lambda_{2}$, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}_{\exp}$ which is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$, where
$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.14)
@@
Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed. However, for a double check, we give its independent proof in Appendix A.2. This proof is straightforward because $L^{(r)}(\cdot)$ is a birth-and-death process, and helpful to intuitively convince (ii) of Theorem 3.1.
@@ 4 Proofs
In this section we prove Theorem 3.1. Compared with the exponential case, a major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$. Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below. This method is called a BAR approach, and has been used in [2, 3, 10]. Recently, this approach is further studied in [11].
@@ 4.1 Palm distribution and BAR
$X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{0},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{0},\end{cases}$
$L^{(r)}_{1}(t)=L(t-)+1(\Delta N_{e}(t)=1),\qquad R^{(r)}_{1}(t)=(R^{(r)}_{e}(t),R^{(r)}_{s}(t-)),$
where $n_{e,1}(t)=\sum_{n=1}^{N_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{0})$ and $n_{e,2}(t)=\sum_{n=1}^{N_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{0})$. We call $X^{(r)}_{1}(t)$ an intermediate state. Note that $X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N_{s}(t)))$ when $\Delta N_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N_{s}(t)=0$.
@@
    Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N_{e}$ is counted at time 00 and just before when $N_{s}$ is counted at time 00. This means that we take a so called arrival first framework. As for the intermediate state, see [11] for its details. We use $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$. This prevents to explicitly refer to the intermediate state $X^{(r)}_{1}(0)$. Of course, the intermediate state is not needed if $N^{(r)}_{e}$ and $N^{(r)}_{s}$ do not have a simultaneous count, but such restriction is avoided by introducing $X^{(r)}_{1}(0)$.
@@
For each integer $k\geq 0$, let $f({\boldsymbol{x}})=x_{1}\wedge(k+1)$ for ${\boldsymbol{x}}\in S$, then obviously $f\in C_{b}^{p}(S)$. Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N_{e}(u)-N_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N_{s}(u)-N_{s}(u-)=1)$, we have
@@
To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$. We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 4.3.

### v5

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study the limit of the stationary distribution of this two-level queue under the scaling in heavy traffic which corresponds to a diffusion approximation. It is shown that this limiting distribution is truncated exponential below the threshold level and exponential above it under suitably chosen system parameters. This result is proved using the so called BAR approach studied in [2, 3, 10, 11]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distributions.
@@ 1 Introduction
This 2-level $GI/G/1$ queue is motivated by an energy saving problem to control the arrival process and service rate. Here, the system is an internet server, and its energy consumption depends on service rate. Then, the problem is to find an optimal threshold to minimize the operating energy of the server and customer’s waiting cost.
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. It may be considered to approximate it by a quasi-birth-and-death process, then it is not hard to numerically compute its performance. However, it is hard to derive it in an analytically tractable form except for special cases such as all the distributions to be exponential. Because of this, a simple approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution. This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
@@
As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue (e.g., see Kingman [7]). Furthermore, Miyazawa [11] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room. Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see Harrison [5]). Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it. This conjecture can be directly verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional. Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [2, 3, 10, 11]. This verification, Theorem 3.1, requires either some limitation on system parameters or for the workload brought by a customer to be exponentially distributed. The proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
@@
This paper is made up by five sections. We first introduce a general framework for a Markov process describing the 2-level $GI/G/1$ queue in Section 2, then provide notations and assumptions for a sequence of the 2-level $GI/G/1$ queues in heavy traffic in Section 3.1. A main result, Theorem 3.1, and the exponentially distributed case, Proposition 3.1, are presented in Section 3.2. In Section 4, we prove Theorem 3.1. Finally, in Section 5, we intuitively discuss two topics. In Section 5.1, we consider a process-limit for the 2-level $GI/G/1$ queue in heavy traffic. In Section 5.2, we intuitively verify the conjecture in Remark 3.1.
@@ 3.1 Assumptions
For the $r$-th model, let $\ell^{(r)}_{0}$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$. We refer to this number as the queue length. Let $t^{(r)}_{e,n}$ be the $n$-th arrival time of a customer, and let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customer. We assume that $T_{e}^{(r)}(n)\equiv t^{(r)}_{e,n}-t^{(r)}_{e,n-1}$ is independent of $\{t^{(r)}_{e,n^{\prime}};0\leq n^{\prime}\leq n-1\}$ and everything else for all $n\geq 1$, where $t^{(r)}_{e,0}=0$. Thus, the inter-arrival times are independent, but they may have two different distributions depending on the queue length. Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are random variables which may have different distributions. As for service, let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customers. We assume that $\{T^{(r)}_{s}(n);n\geq 1\}$ is a sequence of $i.i.d.$ random variable. We denote a random variable subject to the common distribution of $T^{(r)}_{s}(n)$ by $T^{(r)}_{s}$. Let $c^{(r)}_{1}$ and $c^{(r)}_{2}$ be the service speeds when the queue length is not greater than $\ell^{(r)}_{0}$ and otherwise, respectively. Namely, as long as $R^{(r)}_{s}(t)$ is differentiable at time $t$,
@@
Figure 3.1: Sample path of the queue length process
@@
and denote the counting processes of the arrivals and service completions by $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$, respectively. Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov. Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with the time-shift operator $\Psi$. We assume the following assumptions.
@@
    There are constants $b_{1}\in\mathbb{R}$ and $b_{2}>0$ such that $c^{(r)}_{i}\mu^{(r)}-\lambda^{(r)}_{i}=r\mu^{(r)}b_{i}+o(r)$ as $r\downarrow 0$.
@@
    There is a constant $\ell_{0}$ such that $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$, and $\ell^{(r)}_{0}\geq 2$.
@@
Here, we write $f(r)=o(g(r))$ as $r\downarrow 0$ for positive valued functions $f,g$ of $r$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$. Similarly, $f(r)=O(g(r))$ if $\limsup_{r\downarrow 0}f(r)/g(r)<\infty$. Note that (3.a) implies that
@@
Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$. Obviously, $\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,0}(t)]$ under the stationary framework. Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}(n)\wedge T_{e,2}(n)]<\infty$. Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)]<\infty$. ∎
@@ 3.2 Main results
We are now ready to present a main result, which will be proved in Section 4. Recall that $L^{(r)}$ is a random variable subject to the stationary distribution of the $r$-th 2-level $GI/G/1$ queue. Define probability distribution $\nu^{(r)}$ under the assumptions (3.a)–(3.e) as
$\nu^{(r)}(B)=\mathbb{P}(rL^{(r)}\in B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.1)
We refer to this $\nu^{(r)}$ as a scaled stationary distribution. We decompose $\nu^{(r)}$ into measures $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$ as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
@@
$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}>\ell^{(r)}_{0}).$
@@
For each $r\in(0,1]$, define $\widetilde{\nu}_{i|cd}$ as the probability measures on ${\mathcal{B}}(\mathbb{R}_{+})$ which have the density functions $h_{i|cd}$ for $i=1,2$ such that
@@
Conversely, if (3.12) holds, then we have the weak convergence (3.11) from (i) because $\widetilde{A}_{i}>0$. Thus, (3.12) is equivalent to (3.11). We conjecture that (3.11) holds without any condition in (ii). There are two supports for this conjecture. One is from a process limit for the 2-level $GI/G/1$ queue in heavy traffic. Another is from the intuitive verification of (3.12). They are detailed in Sections 5.1 and 5.2, repectively.
@@
Let consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed. We call this model the $r$-th 2-level $M/M/1$ queue.
@@
Under the settings of Theorem 3.1, the scaled stationary distribution $\nu^{(r)}_{\exp}$ of the $r$-th 2-level $M/M/1$ queue weakly converges to $\widetilde{\nu}_{\exp}$ as $r\downarrow 0$, where $\widetilde{\nu}_{\exp}$ is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$ in which
$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{\beta_{2}\ell_{0}}{1+\beta_{2}\ell_{0}},&b_{1}=0,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.15)
@@
Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed. However, for a double check, we give its direct proof in Appendix A.2, which is helpful to intuitively convince (ii) of Theorem 3.1. This proof is straightforward, but one must be careful because $L^{(r)}(\cdot)$ itself is not a continuous-time Markov chain even under the exponential assumptions (see Appendix A.2 for its reason).
@@ 4 Proofs
In this section we prove Theorem 3.1. A major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$. Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below. This method is called a BAR approach, and has been used in [2, 3, 10]. Recently, this approach is further studied in [11].
@@ 4.1 Palm distribution and BAR
$X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N^{(r)}_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{0},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{0},\end{cases}$
$L^{(r)}_{1}(t)=L(t-)+1(\Delta N^{(r)}_{e}(t)=1),\qquad R^{(r)}_{1}(t)=(R^{(r)}_{e}(t),R^{(r)}_{s}(t-)),$
where $n_{e,1}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{0})$ and $n_{e,2}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{0})$. We call $X^{(r)}_{1}(t)$ an intermediate state. Note that $X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N^{(r)}_{s}(t)))$ when $\Delta N^{(r)}_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N^{(r)}_{s}(t)=0$.
@@
    Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N^{(r)}_{e}$ is counted at time 00 and just before when $N^{(r)}_{s}$ is counted at time 00. This means that we take a so called arrival first framework. As for the intermediate state, see [11] for its details. We use $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$. This prevents to explicitly refer to the intermediate state $X^{(r)}_{1}(0)$. Of course, the intermediate state is not needed if $N^{(r)}_{e}$ and $N^{(r)}_{s}$ do not have a simultaneous count, but such restriction is avoided by introducing $X^{(r)}_{1}(0)$.
@@
For each integer $k\geq 0$, let $f({\boldsymbol{x}})=x_{1}\wedge(k+1)$ for ${\boldsymbol{x}}\in S$, then obviously $f\in C_{b}^{p}(S)$. Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N^{(r)}_{e}(u)-N^{(r)}_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N^{(r)}_{s}(u)-N^{(r)}_{s}(u-)=1)$, we have
@@
To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$, which is possible by Lemma 4.1. We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 4.3.